OpenStudy (anonymous):

probability of getting 4 heads in 4 flips? (fraction)

OpenStudy (anonymous):

\[\left( \frac{ 1 }{ 2 } \right)^4=?\]

OpenStudy (anonymous):

1/16?

OpenStudy (anonymous):

1/2 ^ 1/2 ^ 1/2 ^ 1/2

OpenStudy (anonymous):

so yes